OpenStudy (anonymous):

Consider the function h(x) = 3x^3-7x^2+6x-8 What type of function is this? I think it's a polynomial for sure.. um.. could it be exponential, radical, rational, linear?

jimthompson5910 (jim_thompson5910):

This is a cubic polynomial function.

OpenStudy (anonymous):

Polynomial, cubic, Rational

OpenStudy (anonymous):

what makes it rational?

OpenStudy (anonymous):

domain and range both all real right?

OpenStudy (anonymous):

p(x) = 3x^3-7x^2+6x-8 q(x) = 1 A rational function is p(x)/q(x)

OpenStudy (anonymous):

pk that makes as much sense as your name does.

OpenStudy (anonymous):

It's polynomial. For the exponential it means that the exponent is variable like this: \[4^{2x}\]

OpenStudy (anonymous):

:D thanks Raheen!

OpenStudy (anonymous):

But what makes it rational?

jimthompson5910 (jim_thompson5910):

Any function h(x) in the form \[h(x)=\frac{f(x)}{g(x)}\] is a rational function where f(x) and g(x) are polynomial functions. In your case, if f(x)=3x^3-7x^2+6x-8 and g(x)=1, then \[h(x)=\frac{f(x)}{g(x)}=\frac{3x^3-7x^2+6x-8}{1}=3x^3-7x^2+6x-8\], so it's technically a rational function.

OpenStudy (anonymous):

but then any quadratic could be rational as well..?

jimthompson5910 (jim_thompson5910):

Technically, yes, any quadratic is rational. I would stick to my first answer in saying that it's a cubic polynomial function since that's the most accurate description.

OpenStudy (anonymous):

is it also a quadratic?

jimthompson5910 (jim_thompson5910):

No, quadratic functions are of degree 2 (ie max exponent is 2), but the max exponent here is 3.

OpenStudy (anonymous):

I see. :)